# Stratified Hazard Sampling: Minimal-Variance Event Scheduling for CTMC/DTMC Discrete Diffusion and Flow Models

### v1

@@ 6 Conclusion
We proposed Stratified Hazard Sampling (SHS), a drop-in, hyperparameter-free variance-reduction rule for step-based inference in CTMC/DTMC discrete diffusion and flow models with a (stay vs. replace) structure. By stratifying cumulative hazard/jump mass with a single random phase per position, SHS preserves the expected number of edits while achieving minimal possible jump-count variance ($\leq 1/4$) without changing destination sampling.
Across quantized MNIST and a lightweight UDLM text-generation protocol, SHS improves stability and sample quality in the few-step / low-NFE regime. Future work will scale these findings to larger language/code benchmarks and evaluate safety-oriented extensions such as lexically constrained sampling.
Appendix A Full Proof of Proposition 4.1 (Unbiasedness)
We restate the proposition in the scalar form. Fix $S\geq 0$ and draw $\theta\sim\mathrm{Uniform}(0,1)$. Let $I=\lfloor S\rfloor$ and $f=\{S\}=S-I\in[0,1)$. Define
$J\;=\;I+\mathbf{1}[\theta<f].$
@@
$\mathbb{E}[J]\;=\;I+\mathbb{P}(\theta<f)\;=\;I+f\;=\;S,$
since $\theta$ is uniform on $[0,1]$. This proves Proposition 4.1.
Appendix B Full Proof of Proposition 4.2 (Minimal Variance)
Using the notation from Appendix A, we have $J=I+B$ where $B=\mathbf{1}[\theta<f]\sim\mathrm{Bernoulli}(f)$. Since $I$ is deterministic,
$\mathrm{Var}(J)=\mathrm{Var}(I+B)=\mathrm{Var}(B)=f(1-f)\leq\frac{1}{4},$
@@
Minimality among two-point unbiased estimators.
Any integer-valued random variable $\widetilde{J}$ supported on $\{I,I+1\}$ is determined by $p=\mathbb{P}(\widetilde{J}=I+1)$. Unbiasedness $\mathbb{E}[\widetilde{J}]=S=I+f$ forces $p=f$. Therefore $\mathrm{Var}(\widetilde{J})=p(1-p)=f(1-f)$ for any such unbiased estimator, and SHS attains this minimum. This proves Proposition 4.2.
Appendix C Additional Analysis of Stratified Hazard Sampling
@@
C.1 SHS as a systematic coupling of Bernoulli trials
@@
SHS instead draws a single phase $\theta\sim\mathrm{Uniform}(0,1)$ and defines a coupled indicator sequence by
$B_{k}^{\mathrm{shs}}\;=\;\mathbf{1}\!\left[\;\lfloor S_{k}\rfloor+\mathbf{1}[\theta<\{S_{k}\}]\;<\;\lfloor S_{k}+p_{k}\rfloor+\mathbf{1}[\theta<\{S_{k}+p_{k}\}]\;\right],$ (18)
which is equivalent to triggering a change whenever the cumulative sum crosses the next unit boundary $\theta+m$. Summing (18) over $k$ yields the randomized-rounding form $J^{\mathrm{shs}}=\lfloor S_{n}\rfloor+\mathbf{1}[\theta<\{S_{n}\}]$ (cf. Eq. (9)).
C.2 Event-time stratification in cumulative hazard space
@@
C.3 Scope: inference-time variance reduction vs. exact simulation
The results in Propositions 4.14.2 are unconditional statements about randomized rounding for a fixed mass $S$ (or a fixed sequence $\{p_{k}\}$). In a full generative model, the per-step masses $p_{ik}(x)$ (or rates $\lambda_{i}(t,x)$) are typically state-dependent. Because SHS couples change decisions over time through the phase $\theta_{i}$, it introduces additional memory beyond the visible state $x$. Accordingly, SHS should be viewed as an inference-time variance reduction integrator for CTMC/DTMC samplers rather than as an exact simulator of the underlying Markov process.
@@
Appendix D Safety Evaluation Details
This appendix details the evaluation protocol used to support the claim in Section 4.2 that our phase allocation reduces the distributional mismatch to the conditional target $p_{1}(\cdot\mid A)$ compared to naive masking.
D.1 Setup and Compared Samplers
Lexical constraint.
Let $\mathcal{V}$ be the vocabulary and $\mathcal{B}\subset\mathcal{V}$ a blacklist. The safe set is
$\mathcal{A}\;=\;\{x\in\mathcal{V}^{N}:\forall i,\;x_{i}\notin\mathcal{B}\}.$ (20)
The ideal constrained target distribution is the conditional terminal distribution $p_{1}(x\mid\mathcal{A})\propto p_{1}(x)\,\mathbf{1}[x\in\mathcal{A}]$ (cf. Eq. (5) in the main text). We compare:
  • Unconstrained: baseline sampler without any censorship.
  • Mask-only (naive): at each jump, sample from the masked destination $q^{\mathrm{MASK}}_{t,i}(v\mid x)\propto q_{t,i}(v\mid x)\mathbf{1}[v\notin\mathcal{B}]$, while keeping the original escape rates $\lambda_{i}(t,x)$ unchanged.
  • Mask + phase allocation (ours): same masked destination as above, but allocate SHS phases $\{\theta_{i}\}$ by assigning smaller pre-sampled phases to higher-risk positions (Section 4.2, Eq.( 14)), while keeping the original escape rates $\lambda_{i}(t,x)$ unchanged.
  • Rejection reference (optional): samples from the unconstrained sampler followed by rejection of sequences not in $\mathcal{A}$, used as an empirical proxy for $p_{1}(\cdot\mid\mathcal{A})$ (Section D.4).
Why the reference is meaningful.
Since $p_{1}(x\mid\mathcal{A})\propto p_{1}(x)\mathbf{1}[x\in\mathcal{A}]$, acceptance–rejection from $p_{1}$ yields exact conditional samples when the base sampler is exact, and provides a practical approximation otherwise. We report the acceptance rate to contextualize how challenging the constraint is.
D.2 Risk Score and Safety-Critical Edits
Our method accelerates positions that are likely to be affected by censorship. To make this measurable, we define a pre-mask risk score from the model’s unfiltered destination distribution.
Pre-mask forbidden mass.
Given the current state $x$ at time $t$, define for each position $i$:
$r_{i}(t,x)\;=\;\sum_{v\in\mathcal{B}}q_{t,i}(v\mid x),$ (21)
where $q_{t,i}(\cdot\mid x)$ is computed before applying the blacklist filter. Intuitively, $r_{i}(t,x)$ estimates how much the unconstrained dynamics “wants” to place a forbidden token at position $i$ at time $t$.
Safety-critical (censorship-active) jump.
We call a jump at position $i$ and time $t$ safety-critical if
$r_{i}(t,X_{t^{-}})\;\geq\;\tau,$ (22)
for a fixed threshold $\tau\in(0,1)$. This focuses the analysis on edits where masking materially changes the local transition. In all safety experiments, we report $\tau$ and include a sensitivity check over a small grid (e.g., $\tau\in\{0.01,0.05,0.1\}$).
Remark (robust variant).
If the initial distribution $p_{0}$ can contain blacklisted tokens, we additionally mark an edit as safety-critical whenever $X_{t^{-}}^{i}\in\mathcal{B}$, since such a token must eventually be removed to satisfy $\mathcal{A}$.
D.3 Time-to-First-Safe-Edit CDF
To quantify the “late masking” failure mode, we measure how early the sampler performs a safety-critical edit under each censorship method.
Definition.
Let $\{t_{k}\}_{k=0}^{n}$ be the discretization times ($t_{k}=kh$). For a sampled trajectory $(X_{t_{k}})_{k=0}^{n}$, define the time-to-first-safe-edit
[…]
with the convention $T_{\mathrm{edit}}=1$ if no such edit occurs.
Empirical CDF.
Given $M$ independent runs, we estimate
$\widehat{F}_{\mathrm{edit}}(t)\;=\;\frac{1}{M}\sum_{m=1}^{M}\mathbf{1}\!\left[T^{(m)}_{\mathrm{edit}}\leq t\right].$ (24)
We plot $\widehat{F}_{\mathrm{edit}}(t)$ for Mask-only vs. Ours. A left-shift (larger CDF at the same $t$) indicates that safety-critical positions are resolved earlier, consistent with our claim that phase allocation mitigates late masking.
Optional complementary curves.
We also report: (i) time-to-last-safe-edit $T_{\mathrm{last}}$ (the last time a safety-critical jump occurs), and (ii) the unsafe survival curve $t\mapsto\mathbb{P}(\exists i:r_{i}(t,X_{t})\geq\tau)$, both of which directly diagnose “late” censorship activity.
D.4 Approximating $p_{1}(\cdot\mid\mathcal{A})$ by Rejection
When feasible, we build a reference set of samples approximating $p_{1}(\cdot\mid\mathcal{A})$ via rejection:
  1. 1.
    Run the Unconstrained sampler to produce terminal samples $x^{(1)},x^{(2)},\dots$.
  2. 2.
    Accept samples with $x^{(m)}\in\mathcal{A}$ until collecting $M_{\mathrm{ref}}$ accepted samples.
We denote the resulting empirical reference distribution by $\widehat{p}_{\mathrm{ref}}$. We report the acceptance rate $\hat{\alpha}=M_{\mathrm{ref}}/M_{\mathrm{drawn}}$.
Practical note.
If $\hat{\alpha}$ is too small for a given blacklist, we (i) shorten the generated length, (ii) use a milder blacklist, or (iii) treat the reference as approximate and rely on the two-sample metrics in Section D.5 without claiming exactness.
D.5 Terminal Distribution Shift Induced by Censorship
We quantify how much censorship distorts the terminal distribution relative to the conditional target. Let $\mathcal{S}_{\mathrm{ref}}=\{x^{(m)}_{\mathrm{ref}}\}_{m=1}^{M_{\mathrm{ref}}}$ be the rejection reference and $\mathcal{S}_{\mathrm{meth}}=\{x^{(m)}_{\mathrm{meth}}\}_{m=1}^{M}$ be samples produced by a censorship method (Mask-only or Ours). We report multiple complementary discrepancy measures.
(1) Per-position token marginal JSD.
For each position $i$, define empirical marginals
$\hat{p}^{\mathrm{ref}}_{i}(v)=\frac{1}{M_{\mathrm{ref}}}\sum_{m=1}^{M_{\mathrm{ref}}}\mathbf{1}[x^{(m)}_{\mathrm{ref},i}=v],\quad\hat{p}^{\mathrm{meth}}_{i}(v)=\frac{1}{M}\sum_{m=1}^{M}\mathbf{1}[x^{(m)}_{\mathrm{meth},i}=v].$ (25)
We compute the Jensen–Shannon divergence
$\mathrm{JSD}_{i}=\mathrm{JSD}\!\left(\hat{p}^{\mathrm{meth}}_{i}\,\|\,\hat{p}^{\mathrm{ref}}_{i}\right),\quad\mathrm{JSD}(p\|q)=\frac{1}{2}\mathrm{KL}\!\left(p\Big\|\frac{p+q}{2}\right)+\frac{1}{2}\mathrm{KL}\!\left(q\Big\|\frac{p+q}{2}\right),$ (26)
and report the average $\frac{1}{N}\sum_{i=1}^{N}\mathrm{JSD}_{i}$.
(2) Sequence-level feature divergence.
To capture higher-order effects beyond position-wise marginals, we additionally compute a divergence on lightweight sequence features $\phi(x)$, such as: (i) unigram/bigram count vectors, or (ii) a fixed pretrained sentence embedding. For count-based features, we compute JSD between normalized feature histograms. For embeddings, we report a kernel two-sample statistic (MMD) or a linear-probe classification accuracy (two-sample test) as a proxy for distributional shift.
Interpretation.
Smaller discrepancy to $\hat{p}_{\mathrm{ref}}$ indicates closer approximation to $p_{1}(\cdot\mid A)$. We expect Mask-only to incur larger shift because it only filters one-step destinations (Eq. (8)) and does not capture the lookahead-dependent scheduling effects implied by the Doob $h$-transform (Eq. (7)). In contrast, our phase allocation prioritizes early edits at high-risk positions, reducing late masking artifacts and the resulting terminal distribution shift.

### v2

@@ 6 Conclusion
We proposed Stratified Hazard Sampling (SHS), a drop-in, hyperparameter-free variance-reduction rule for step-based inference in CTMC/DTMC discrete diffusion and flow models with a (stay vs. replace) structure. By stratifying cumulative hazard/jump mass with a single random phase per position, SHS preserves the expected number of edits while achieving minimal possible jump-count variance ($\leq 1/4$) at virtually no computational cost, without changing destination sampling.
Experiments on UDLM, GIDD, and a 3B-parameter uniform diffusion model (von Rütte et al., 2025b) show consistent sample-quality gains, especially in the few-step regime, and improved robustness under increasingly severe vocabulary truncation. MNIST diagnostics further confirm the intended regularization of event timing (Appendix A). Future work will extend experiments to larger code benchmarks and discrete flow matching models.
Acknowledgments and Disclosure of Funding
Appendix A MNIST test
A.1 Setup
Our goal is to isolate the effect of event-time sampling on a continuous-time discrete generative model. We fix the pretrained predictor and vary only the sampling procedure: (i) the standard sampler, and (ii) our Stratified Hazard Sampling (SHS).
Few-step regime.
To stress-test trajectory instability, we vary the number of discretization steps $\text{NFE}\in\{256,64,16,8\}$ while keeping the same model and schedule.
(a) Standard sampling
(b) SHS
Figure 4: Hazard-space jump locations. Histograms of $S_{k}$ (the cumulative hazard at the $k$-th jump; shown as cumulative_prob on the x-axis in the plots). Standard sampling exhibits Erlang/Gamma-shaped variability, while SHS produces stratified, bounded-support locations, demonstrating reduced trajectory randomness.
A.2 Jump Timing Diagnostics in Hazard Space
$S(t)=\int_{0}^{t}\lambda(s)\,ds.$ (8)
We record the hazard-space jump locations $\{S_{k}\}_{k\geq 1}$, where $S_{k}$ is the cumulative hazard value at the $k$-th jump.
Standard baseline: Erlang/Gamma-shaped locations.
Under the standard NHPP view, inter-event increments in hazard space satisfy $\Delta S_{k}\sim\mathrm{Exp}(1)$ i.i.d., so the $k$-th jump location becomes
$S_{k}=\sum_{i=1}^{k}\Delta S_{i}\sim\mathrm{Gamma}(k,1),$ (9)
which yields an Erlang-shaped distribution as $k$ increases. Empirically, we observe broad, heavy-tailed variability in early jump locations under standard sampling, consistent with the “under-/over-edit” instability in the noise-start setting.
SHS: stratified (bounded-support) jump locations.
A.3 Quantized MNIST Reconstructions: Eliminating Under-Edit
We next visualize the practical effect of trajectory stabilization on a simple discrete generation task. Starting from a noise initialization, we run the sampler for $\text{NFE}\in\{256,64,16,8\}$ discretization steps and compare the resulting reconstructions.
Observation.
Standard sampling often exhibits an under-edit failure mode in the few-step regime: some positions experience too few jumps, leaving noticeable $x_{0}$-derived noise even late in sampling. In contrast, SHS produces visibly cleaner samples at early stages and substantially reduces residual noise in low-NFE runs. Qualitatively, SHS maintains legible digit structure even at $\text{NFE}=16$ and $\text{NFE}=8$, where Standard outputs remain dominated by unresolved noise.
$\text{NFE}=256$
$\text{NFE}=64$
$\text{NFE}=16$
$\text{NFE}=8$
(a) Standard sampling
$\text{NFE}=256$
$\text{NFE}=64$
$\text{NFE}=16$
$\text{NFE}=8$
(b) SHS
Figure 5: Quantized MNIST reconstructions under varying step budgets. Standard sampling vs. SHS at $\text{NFE}\in\{256,64,16,8\}$ discretization steps. SHS reduces the under-edit regime and yields cleaner early-stage reconstructions.
Appendix B Full Proof of Proposition 1 (Unbiasedness)
Fix $S\geq 0$ and draw $\theta\sim\mathrm{Uniform}(0,1)$. Let $I=\lfloor S\rfloor$ and $f=S-I\in[0,1)$. Define
$J\;=\;I+\mathbf{1}[\theta<f].$ (10)
@@ Appendix B Full Proof of Proposition 1 (Unbiasedness)
$\mathbb{E}[J]\;=\;I+\mathbb{P}(\theta<f)\;=\;I+f\;=\;S,$ (11)
since $\theta$ is uniform on $[0,1]$. This proves Proposition 1.
Appendix C Full Proof of Proposition 2 (Minimal Variance)
Using the notation from Appendix B, we have $J=I+B$ where $B=\mathbf{1}[\theta<f]\sim\mathrm{Bernoulli}(f)$. Since $I$ is deterministic,
$\mathrm{Var}(J)=\mathrm{Var}(I+B)=\mathrm{Var}(B)=f(1-f)\leq\frac{1}{4},$ (12)
@@ Appendix C Full Proof of Proposition 2 (Minimal Variance)
Finally, any integer-valued random variable $\widetilde{J}$ supported on $\{I,I+1\}$ is determined by $p=\mathbb{P}(\widetilde{J}=I+1)$. Unbiasedness $\mathbb{E}[\widetilde{J}]=S=I+f$ forces $p=f$. Therefore $\mathrm{Var}(\widetilde{J})=p(1-p)=f(1-f)$ for any such unbiased estimator, and SHS attains this minimum. This proves Proposition 2.
Appendix D Additional Analysis of Stratified Hazard Sampling
@@
D.1 SHS as a systematic coupling of Bernoulli trials
@@
SHS instead draws a single phase $\theta\sim\mathrm{Uniform}(0,1)$. Write $S_{k}=I_{k}+f_{k}$ with $I_{k}:=\lfloor S_{k}\rfloor$ and $f_{k}:=S_{k}-I_{k}\in[0,1)$, and note that $S_{k+1}=S_{k}+p_{k}$ (so $S_{k+1}=I_{k+1}+f_{k+1}$ with $I_{k+1}:=\lfloor S_{k+1}\rfloor$ and $f_{k+1}:=S_{k+1}-I_{k+1}$). We define a coupled indicator sequence by
$B_{k}^{\mathrm{shs}}\;=\;\mathbf{1}\!\left[\;I_{k}+\mathbf{1}[\theta<f_{k}]\;<\;I_{k+1}+\mathbf{1}[\theta<f_{k+1}]\;\right],$ (13)
which is equivalent to triggering a change whenever the cumulative sum crosses the next unit boundary $\theta+m$. Summing equation 13 over $k$ yields the randomized-rounding form $J^{\mathrm{shs}}=\lfloor S_{n}\rfloor+\mathbf{1}[\theta<f_{n}]$ (cf. equation 6).
D.2 Event-time stratification in cumulative hazard space
@@
D.3 Scope: inference-time variance reduction vs. exact simulation
The results in Propositions 12 are unconditional statements about randomized rounding for a fixed mass $S$ (or a fixed sequence $\{p_{k}\}$). In a full generative model, the per-step masses $p_{ik}(x)$ (or rates $\lambda_{i}(t,x)$) are typically state-dependent. Because SHS couples change decisions over time through the phase $\theta_{i}$, it introduces additional memory beyond the visible state $x$. Accordingly, SHS should be viewed as an inference-time variance reduction integrator for CTMC/DTMC samplers rather than as an exact simulator of the underlying Markov process.
@@
Composability with destination-side modifications.
SHS changes only the event scheduling (stay vs. replace decisions) by coupling per-step Bernoulli trials through a single phase, and leaves the per-jump destination sampling unchanged (Algorithm 1). Therefore, SHS can be composed with destination-side modifications such as vocabulary truncation / blacklist filtering by simply replacing $q_{t,i}$ (or $q_{k,i}$) with the filtered destination $q^{(\rho)}$ in equation 5, while keeping the same change masses $p_{ik}$ (or escape rates $\lambda_{i}$).
D.4 Absorbing-state (mask-start) implies single-jump trajectories
This appendix formalizes why the jump-count variance discussed in Section 2.1 is a multi-edit phenomenon and is absent under the standard absorbing-state (mask-start) setting.
Setup.
Let $V$ be the vocabulary and let $m=\texttt{[MASK]}$ be a dedicated sentinel token. Define the extended alphabet $\bar{V}:=V\cup\{m\}$, and consider step-based sampling on a grid $\{t_{k}\}_{k=0}^{n}$. A state is $x\in\bar{V}^{N}$. At each step $k$ and position $i$, assume the sampler admits a (stay vs. replace) decomposition as in Section 2.1, with a change mass $p_{ik}(x)\in[0,1]$ and a destination distribution $q_{k,i}(\cdot\mid x)\in\Delta(V)$.
Unmask-only (absorbing) assumption.
In the absorbing-state (mask-start) sampler, we assume unmask-only dynamics: once a position leaves $m$, it never changes again. Concretely, for any $x\in\bar{V}^{N}$,
$p_{ik}(x)=0\quad\text{whenever }x_{i}\in V.$ (15)
When $x_{i}=m$, the update at step $k$ draws $B_{ik}\sim\mathrm{Bernoulli}(p_{ik}(x))$ and, if $B_{ik}=1$, sets $x^{\prime}_{i}\sim q_{k,i}(\cdot\mid x)$ (thus $x^{\prime}_{i}\in V$); otherwise it keeps $x^{\prime}_{i}=m$.
Define the (non-trivial) edit indicator and total edit count
$B_{ik}:=1[x^{\prime}_{i}\neq x_{i}],\qquad J_{i}:=\sum_{k=0}^{n-1}B_{ik}.$ (16)
Proposition 3 (Single-edit property under absorbing-state dynamics).
Under equation 15, each position can undergo at most one edit: for every $i$, $J_{i}\in\{0,1\}$ almost surely.
Proof.
If $J_{i}=0$ there is nothing to prove. Otherwise, let $k^{\star}$ be the first step where $B_{ik^{\star}}=1$. By construction, $x_{i}$ changes only when $x_{i}=m$, hence $x^{(k^{\star})}_{i}=m$ and after the update we have $x^{(k^{\star}+1)}_{i}\in V$. Then equation 15 implies that for all later steps $\ell>k^{\star}$, $p_{i\ell}(x^{(\ell)})=0$ and thus $B_{i\ell}=0$. Therefore exactly one edit can occur, so $J_{i}=1$. ∎
Variance implication.
Since $J_{i}\in\{0,1\}$, it is a Bernoulli random variable and
$\mathrm{Var}(J_{i})\leq\tfrac{1}{4}.$ (17)
Moreover, many mask-start implementations explicitly force any remaining [MASK] tokens to be resolved by the terminal time (e.g., by setting $p_{i,n-1}(x)\equiv 1$ for $x_{i}=m$ at the final step), in which case $J_{i}=1$ deterministically and $\mathrm{Var}(J_{i})=0$.
Interpretation (why SHS is mainly needed for uniform-noise start).
Under mask-start, the state explicitly reveals whether a position has been generated via the sentinel $m=\texttt{[MASK]}$, and the unmask-only constraint equation 15 enforces a single-edit trajectory per position. In contrast, uniform-noise start typically requires multiple self-correction edits per position, where sampler randomness in the number and timing of edits becomes substantial (Section 2.1), motivating SHS as a minimal-variance event scheduler.
Appendix E Blacklists: conditioning lens and mass-preserving filtering
This appendix collects background and technical details for the random-blacklist experiments in Section 5.2. We emphasize that our main paper studies SHS as an inference-time variance reduction rule (Appendix D, Section D.3), and we use blacklists primarily as a robustness stress test rather than as an exact conditional sampler.
E.1 Conditioning viewpoint and Doob $h$-transform (background)
Let $\mathcal{B}\subset\mathcal{V}$ be a blacklist and define the allowed event $\mathcal{A}=\{x\in\mathcal{V}^{N}:\forall i,\ x_{i}\notin\mathcal{B}\}$. If the goal is the conditional terminal distribution,
$p_{1}(x\mid\mathcal{A})\;=\;\frac{p_{1}(x)\,\mathbf{1}[x\in\mathcal{A}]}{\mathbb{P}_{p_{1}}(X_{1}\in\mathcal{A})},$ (18)
then conditioning generally changes the dynamics at intermediate times.
CTMC Doob transform.
For a time-inhomogeneous CTMC with generator $Q_{t}$, define the survival function
$h_{t}(x)\;=\;\mathbb{P}(X_{1}\in\mathcal{A}\mid X_{t}=x).$ (19)
A classical result (Doob’s $h$-transform) gives the conditioned generator
with diagonal entries chosen so rows sum to zero. Crucially, equation 20 reweights not only which transitions are taken but also their effective intensities, through the lookahead ratio $h_{t}(y)/h_{t}(x)$.
DTMC analogue.
An analogous $h$-transform exists for discrete-time Markov chains and similarly reweights the transition kernel by a lookahead ratio. The key point is the same: the correct conditional dynamics depends on future survival.
Why destination-only masking is not exact conditioning.
A common blacklist heuristic modifies only the destination distribution via renormalization (cf. equation 5):
$q^{(\rho)}_{t,i}(v\mid x)\ \propto\ q_{t,i}(v\mid x)\,\mathbf{1}[v\notin\mathcal{B}_{\rho}],$ (21)
while leaving the base escape rates $\lambda_{i}(t,x)$ (or DTMC stay-vs.-replace masses $p_{ik}(x)$) unchanged. This generally defines a different dynamics than the Doob-transformed one equation 20, since the latter incorporates the global lookahead $h_{t}$ in both destinations and intensities. In this paper, we do not attempt to estimate $h_{t}$; we use this viewpoint only as a diagnostic lens.
E.2 Mass-preserving destination filtering used in our experiments
In Section 5.2, we use random vocabulary truncation rather than explicit token-removal conditioning. Fix $\rho$ and let $\mathcal{V}_{\rho}=\mathcal{V}\setminus\mathcal{B}_{\rho}$. We start from $X_{0}\sim\mathrm{Unif}(\mathcal{V}_{\rho})^{N}$ (Eq. equation 4) and enforce the blacklist by renormalizing only the destination distribution (Eq. equation 5), keeping the change mass unchanged.
DTMC kernel view.
Let $P_{k,i}(\cdot\mid x)$ be the model’s per-position categorical kernel, and let $p_{ik}(x)=1-P_{k,i}(x_{i}\mid x)$ and $q_{k,i}(\cdot\mid x)$ be the stay-vs.-replace decomposition (Eq. equation 2). Define the filtered destination
$q^{(\rho)}_{k,i}(v\mid x)\;=\;\frac{q_{k,i}(v\mid x)\,\mathbf{1}[v\in\mathcal{V}_{\rho}]}{\sum_{u\in\mathcal{V}_{\rho}}q_{k,i}(u\mid x)}.$ (22)
Then the filtered kernel is
$P^{(\rho)}_{k,i}(x_{i}\mid x)=P_{k,i}(x_{i}\mid x),\qquad P^{(\rho)}_{k,i}(v\mid x)=p_{ik}(x)\,q^{(\rho)}_{k,i}(v\mid x)\quad(v\neq x_{i}),$ (23)
so the total probability of changing the token remains exactly $p_{ik}(x)$.
CTMC generator view.
Similarly, for a CTMC parameterization $Q_{t}(x^{(i\leftarrow v)}\mid x)=\lambda_{i}(t,x)q_{t,i}(v\mid x)$, we keep $\lambda_{i}(t,x)$ unchanged and replace $q_{t,i}$ by $q^{(\rho)}_{t,i}$ in Eq. equation 5, redistributing rate mass only over the allowed vocabulary.
Relation to “compensating” for removed mass.
An equivalent implementation of sampling $v\sim q^{(\rho)}_{t,i}(\cdot\mid x)$ is rejection sampling: draw $v\sim q_{t,i}(\cdot\mid x)$ repeatedly until $v\in\mathcal{V}_{\rho}$. This can be interpreted as “compensating” for the removed blacklist mass by renormalization; it preserves the effective change probability once a change event is triggered.
E.3 A simple SHS property relevant to blacklist robustness
The blacklist experiments typically become more brittle as $\rho$ increases because generation must succeed under a smaller feasible action space. In this regime, under-editing tails (positions that receive too few edits) are especially harmful. The following elementary proposition formalizes an optimal suppression of the zero-edit event under a fixed realized cumulative mass.
Proposition 4 (SHS minimizes the probability of zero edits under fixed cumulative mass).
Fix $S\geq 0$ and let $J\in\mathbb{Z}_{\geq 0}$ be any random variable such that $\mathbb{E}[J]=S$. Then
$\mathbb{P}(J=0)\ \geq\ (1-S)_{+}\;:=\;\max\{0,1-S\}.$ (24)
Moreover, SHS attains equality: if $\theta\sim\mathrm{Unif}(0,1)$ and $J_{\mathrm{shs}}=\lfloor S\rfloor+\mathbf{1}[\theta<S-\lfloor S\rfloor]$, then $\mathbb{P}(J_{\mathrm{shs}}=0)=(1-S)_{+}$.
Proof.
Since $J\geq\mathbf{1}[J\geq 1]$, we have $\mathbb{E}[J]\geq\mathbb{P}(J\geq 1)$. Thus $\mathbb{P}(J=0)=1-\mathbb{P}(J\geq 1)\geq 1-\mathbb{E}[J]=1-S$. Truncating at 0 yields $(1-S)_{+}$. For SHS, if $S\geq 1$ then $J_{\mathrm{shs}}\geq 1$ almost surely. If $S<1$ then $J_{\mathrm{shs}}=\mathbf{1}[\theta<S]$ is Bernoulli(S)(S), hence $\mathbb{P}(J_{\mathrm{shs}}=0)=1-S$. ∎
How to read Proposition 4 in this paper.
Proposition 4 is a statement about randomized rounding for a fixed realized mass $S$. In the full model, the per-step masses $p_{ik}(x)$ are state-dependent, so SHS should be viewed as an inference-time variance reduction rule rather than an exact simulator (Appendix D, Section D.3). Nevertheless, the proposition provides a useful lens: among unbiased integer edit-count constructions with the same expected total mass, SHS is optimal in suppressing the “no-edit” tail, which is consistent with the robustness trends observed under increasing blacklist severity.